# AI-Enhanced 7-Point Checklist for Melanoma Detection Using Clinical Knowledge Graphs and Data-Driven Quantification

### v1

@@ 2.2.3 Adaptive Receptive Field Proximity (ARFP)
AUC Inception-combined [kawahara2018seven] 86.3 79.9 78.9 79.0 82.9 79.9 89.2 76.1 81.5
EmbeddingNet[yap2018multimodal] 82.5 74.5 77.7 77.9 71.3 78.5 84.8 76.9 78.0
TripleNet [ge2017skin] 81.2 73.8 76.3 77.6 76.8 76.0 85.1 79.9 78.3
Constrained-CC [wang2021incorporating] 83.4 76.2 81.8 82.6 78.4 81.1 90.5 78.6 81.6
HcCNN [bi2020multi] 85.6 78.3 77.6 81.3 81.9 82.6 89.8 82.7 82.5
GRM-DC [fu2022graph] 87.6 87.5 81.2 83.6 79.0 83.1 90.8 75.4 83.5
Ours 89.1 84.9 79.7 83.8 87.1 81.3 92.1 81.9 85.0
Prec Inception-combined [kawahara2018seven] 65.3 61.6 52.7 57.8 56.5 70.5 63.0 30.8 57.3
EmbeddingNet [yap2018multimodal] 68.3 52.5 58.5 61.5 76.8 70.8 89.5 36.8 64.4
TripleNet [ge2017skin] 61.8 59.6 61.7 54.7 77.7 65.7 90.6 64.3 67.0
Constrained-CC [wang2021incorporating] 54.4 66.1 62.8 43.0 76.1 77.8 67.4 75.7 65.4
HcCNN [bi2020multi] 62.8 62.3 52.4 65.1 81.6 69.6 91.9 50.0 67.0
GRM-DC [fu2022graph] 65.6 48.4 50.4 82.3 73.5 74.9 67.4 100 70.3
Ours 77.7 65.8 57.1 74.5 77.6 75.7 75.0 37.1 67.6
Sens Inception-combined [kawahara2018seven] 61.4 48.4 51.1 59.7 66.0 62.1 77.3 13.3 54.9
EmbeddingNet [yap2018multimodal] 40.6 33.3 51.1 60.5 96.2 64.4 96.3 23.3 58.2
TripleNet [ge2017skin] 46.5 33.3 39.4 61.3 97.9 67.2 90.0 30.0 58.2
Constrained-CC [wang2021incorporating] 65.3 39.8 52.1 58.5 40.6 63.8 73.3 30.0 52.9
HcCNN [bi2020multi] 58.4 40.9 35.1 55.7 95.2 80.2 92.2 20.0 59.7
GRM-DC [fu2022graph] 59.0 77.5 67.0 39.2 21.9 70.1 69.9 3.6 51.0
Ours 50.0 51.6 46.8 58.9 55.7 65.0 72.0 43.3 55.4
Spec Inception-combined [kawahara2018seven] 88.8 90.7 85.7 80.1 81.3 78.9 89.4 97.5 86.6
EmbeddingNet [yap2018multimodal] 93.5 90.7 88.7 82.7 20.8 78.4 52.0 96.7 75.4
TripleNet [ge2017skin] 90.1 93.0 92.4 76.8 23.6 71.6 60.0 98.6 75.8
Constrained-CC [wang2021incorporating] 81.6 93.7 90.4 75.7 95.6 85.3 91.6 99.2 89.1
HcCNN [bi2020multi] 88.1 92.4 90.0 86.3 41.5 71.6 65.3 98.4 79.2
GRM-DC [fu2022graph] 89.5 79.0 80.3 95.8 96.8 78.8 91.0 100 88.9
Ours 95.5 91.7 89.0 90.7 94.1 83.0 94.3 94.0 91.5
2.3 Gradient Diagnostics with Data-Driven Weighting (GD-DDW)
In this module, we first establish a parallel multi-label classifier designed to address the classification of seven attributes on the checklist. As $X$ is the input, $y_{{\textrm{7p}}_{j}}=[y_{{\textrm{7p}}_{1}},y_{{\textrm{7p}}_{2}},...,y_{{\textrm{7p}}_{7}}]\in\{0,1\}^{7}$ is denoted as the ground truth of all seven attributes in the format of a binary indicator vector. In the proposed method, the focal loss function [lin2017focal] was employed with the objective of enhancing the performance on the dataset that exhibiting a serious imbalance.
The focal loss function for each label can be defined as:
$L_{\textrm{7p}}=\sum_{j=1}^{7}L_{{\textrm{7p}}_{j}}$ (8)
$L={L_{\textrm{7p}}}+{\lambda L_{\textrm{mel}}}$ (10)
where $\lambda$ is the weighting parameter that adjusting the contributions of the two losses, specifically, making sure the loss values are in the same order of magnitude.
3 Experiments
3.1 Materials
In this study, we utilized the EDRA public dataset as the major material, purposefully curated for 7PCL studies and annotated by Kawahara et al [kawahara2018seven]. This dataset comprises paired dermoscopic and clinical images sourced from 1011 patients, each image has a maximum resolution of 768 by 512 pixels. Notably, nine patients lacked clinical images, which were substituted with corresponding dermoscopic ones. Throughout our study, we strictly adhered to the dataset’s official specifications, which stipulated a total of 413 training, 203 validation, and 395 test images. Concerning the labels, the 7PCL attributes were incorporated into the analysis, following the official categorization. Notably, only the melanoma category was retained from the original disease categorization (DIAG). To verify the generalizability of our proposed method, we also used two publicly available datasets, ISIC 2017 and ISIC 2018 [codella2018skin], which also have multiple attributes. The visual attributes contained in ISIC 2017 include milia-like cysts, negative network, pigment network, and streaks. A total of 2000 training, 150 validation, and 600 test images. ISIC 2018 adds an additional attribute, globules, to the 2017 version and adds a portion of the data to make the data distribution adjusted to 2594 training, 100 validation, and 1000 test images.
@@ 3.1 Materials
Table 2: Comparative Ablation Study of Key Modules in Our Proposed Method, Evaluated on the EDRA Dataset
@@
Baseline (Clinical) Baseline (Derm) Derm & Clinical +IEDWG +ARFP +GD-DDW Ours DIAG MEL 75.6 78.2 77.8 83.5 86.5 83.2 89.1 PN TYP 71.7 75.8 79.4 84.2 85.5 81.5 84.9 ATP 69.5 66.9 68.5 83.9 82.5 77.2 80.2 STR REG 79.2 84.2 82.6 87.5 86.7 83.2 83.5 IR 58.5 67.9 71.6 79.4 77.2 75.3 79.7 PIG REG 66.7 67.8 71.3 71.0 70.2 71.4 82.3 IR 68.1 72.0 76.3 81.8 81.9 82.2 83.8 RS PRS 58.4 72.9 71.9 81.0 83.8 79.4 87.1 DaG REG 69.6 67.7 68.5 79.0 78.0 75.5 80.9 IR 67.1 71.3 70.4 81.0 80.3 79.5 81.3 BWV PRS 79.6 83.6 83.9 94.3 93.8 91.2 92.1 VS REG 71.5 75.6 79.1 82.8 83.9 84.9 86.1 IR 66.4 64.3 72.0 77.7 75.8 75.2 81.9 Avg 69.4 72.9 74.9 82.1 82.0 80.0 84.1
3.2 Implementation Details
In our study, we ensured the integrity of our results by employing consistent data processing methods and adhering to uniform model training protocols across various ablation comparison experiments. Our chosen deep learning model framework is ResNet-50, leveraging parameters pretrained on ImageNet, a common practice in other related works. The entire architecture is meticulously fine-tuned end-to-end, utilizing the Adam optimizer [zhang2018improved] with a concurrent learning rate set to 0.00001. We predefined the number of epochs for pretraining to 150, with an early-stopping mechanism in place. If the model’s performance on the validation set fails to improve over 50 consecutive epochs, the training process halts, and the best-performing model is saved. The loss functions utilized are based on focal loss to circumvent the result bias that may arise from imbalanced data. The hyperparameters for the different stages outlined in the methodology were selected based on insights gained from the training process and represent approximate ranges.
The initial stage of the process is the pre-processing of the images. To maintain consistency and minimize the impact of variable image formats, we eliminate the black edges from the images and resize them to 224 by 224 pixels, aligning with the requirements of the ResNet-50 backbone.
In the CD-MFM section, features from different channels undergo two stages of fusion using weighted averaging. Initially, a fusion process occurs between clinical and dermoscopy images, followed by a second fusion stage after incorporating topological relations learned by the graph network. The parameter $\delta$ in Eq. 1, which corresponds to the weight of clinical images, is set to 0.4, reflecting the dominance of dermoscopy in providing features related to 7PCL. Additionally, the parameters $\gamma$ series in Eq. 3 are configured to ${1}/{3}$ to mitigate potential biases stemming from assigning varying weights to the same feature across multiple instances.
In constructing the graph network, we consider only a subset of the possible connections between nodes due to the directionality of the multistep connections. Based on the average number of co-existing labels in the training set and the avoidance of isolated points, the connections are ordered according to the edge weights, ensuring that each node has at least one and no more than three edges. The order of connection $k$ is set to 3 due to the natural limitation of the number of nodes in 7PCL.
In the final diagnostic phase, the parameter $\lambda$, utilized in the global optimization of the loss function in Eq. 10, is set to 1, given that both loss values are of a comparable magnitude. All experiments were conducted on one NVIDIA Tesla V100 GPU card (32 GB memory) at the UBC ARC Sockeye using the PyTorch platform. The duration of each training cycle is approximately one hour. Our code is publicly available at https://github.com/Ryan315/7PGD.
3.3 Evaluation Settings
The study employed commonly used performance metrics such as the area under the receiver operating characteristic curve (AUC), sensitivity (Sens), specificity (Spec), and precision (Prec). To showcase the advantages of our method, we analyzed the results from various perspectives.
Initially, our method was compared with the state-of-the-art (SOTA) methods that utilized the same dataset and addressed the same task. These methods included GRM-DC [fu2022graph], which considered intermodal and intercategory connections; 7PCL-Contrained-CC [wang2021incorporating], considering the mutual connections between attributes; HcCNN, focusing on multi-channel modal fusion [bi2020multi]; TripleNet, incorporating knowledge distillation transfer [ge2017skin]; and the EmbeddingNet work proposed by Yap et al [yap2018multimodal]. Specifically, we used the Inception models from the deep learning work of Kawahara et al. [kawahara2018seven] as a baseline for comparison, as they originally published the data and proposed the first multi-modal, multi-label deep learning fusion method.
Next, ablation experiments were conducted to validate and analyze the role of crucial modules proposed in our method. Firstly, we verified the role of multimodal fusion CD-MFM by comparing unimodal and multimodal modalities. Secondly, the directed topology between features and diagnostics IEDWG, and the adjustable multi-order link ARFP used in graph network learning, were considered. Thirdly, our gradient-based prediction structure, GD-DDW, was employed for comparison with commonly used parallel prediction structures.
Then, as the proposed method draws inspiration from clinical algorithms in diagnosing real diseases, a comparison with clinical algorithms was also necessary. The prediction results of all attributes obtained in the first layer of the GD-DDW structure were utilized. Following the rule of summing the assigned scores in the clinic, different thresholds from 1 to 7 for the diagnostic scores were set to diagnose melanoma. This was then compared with the proposed method.
To validate the proposed method, we conducted comparative experiments on two additional publicly available datasets, ISIC 2017&2018. While the lack of information regarding the diagnosis precludes an evaluation of the impact of GD-DDW, the directed multi-order relationships between the attributes can still be validated.
@@ 3.3 Evaluation Settings
Table 3: Comparison of Weights Value between Proposed Method and Traditional Algorithm.
@@
Traditional 2 1 1 1 1 2 2
Proposed 1.47 0.95 0.93 0.92 0.97 1.42 1.35
Table 4: Performance (AUC) of Proposed Method on ISIC Datasets.
@@
ISIC2017 Milia-like cysts 63.2 68.3 66.3
Negative network 77.9 84.1 86.5
Pigment network 87.3 87.4 89.6
Streaks 96.2 93.9 94.1
Avg 81.2 83.4 84.1
ISIC2018 Globules 68.2 77.1 79.3
Milia-like cysts 60.1 68.8 66.7
Negative network 82.3 87.3 87.9
Pigment network 87.8 86.4 88.0
Streaks 84.3 88.7 89.4
Avg 76.5 81.7 82.3
4 Results
4.1 Comparison to State-of-the-Art Methods
Table 1 presented a comprehensive comparison of the classification performance between the proposed method and several state-of-the-art methods on the EDRA dataset. It demonstrated our method surpassed other SOTA methods with an average AUC of 85.0%. Notably, five out of the eight labels, specifically including melanoma, achieved the highest diagnostic results. Moreover, it was noteworthy that GRM-DC, which was also based on graph convolutional networks, was observed to be the methodology achieving the optimal AUC for the remaining two of three labels. This finding reinforced the notion that the interactions between attributes in 7PCL could not only aid in detecting melanoma but also offer insights for detecting each attribute.
As for the other three metrics, our method also demonstrated strong performance in precision (67.6%), particularly excelling in Mel (77.7%). The sensitivity of our method is 55.4%, particularly for VS, which represents an improvement of over 13% compared to other methods. As VS represents the most unbalanced attribute in this dataset (with only 71 positive subjects out of 1011), the resulting enhancement of this attribute highlights the significance of the interactions between the attributes learned by our approach. Additionally, it achieved the highest specificity (91.5%), excelling in Mel (95.5%) and BWV (94.3%).
4.2 Ablation Studies
Table 2 presented the results of the ablation experiments conducted on the various modules. The first comparison was between unimodality and multimodality. Consistent with previous related studies, the results showed that dermoscopy and clinical imaging fusion generally outperformed unimodality. Additionally, dermoscopy performed better than clinical images with 3% improvement, which was reasonable given that dermoscopy naturally provided more detailed features, especially in the context of 7PCL.
Regarding the impact of different modules on the results, both IEDWG and ARFD contributed to an average AUC improvement of 8%. Particularly, ARFD showed the highest improvement in melanoma’s results, increasing from 77.8% to 86.5%. This underscores the effectiveness of multi-order relationships between attributes in providing diagnostic information. Additionally, the gradient diagnostic structure, designed to mimic clinical diagnostic approaches, enhanced the average AUC by 5%. The integration of these three modules collectively resulted in a notable enhancement in model efficacy, as evidenced by the average AUC reaching 84.1%.
4.3 Comparison to Traditional 7PCL Algorithm
As shown in Table 3, the traditional algorithm employs approximate categorization based on multivariate analysis [argenziano1998epiluminescence], assigning weights of 2 or 1 to major and minor attributes, respectively. In contrast, the method proposed in this study provides more precise weights, ranging from 1.3 to 1.5 for major attributes and from 0.9 to 1 for minor attributes, based on predicted scores for each category. Fig. 6 illustrates the discrepancy between the proposed method and the traditional 7PCL algorithm, based on identical forecasting outcomes for each attribute. The integration of predicted scores and data-driven weighting modules significantly enhances the prediction of melanoma. The predicted AUC improves from 72% to 89%, representing a substantial increase of 17%. Additionally, the ROC curve of the proposed method provides more accurate information regarding paired specificity and sensitivity, assisting dermatologists in more effective selection and analysis. While traditional methods offer initial guidance concisely, our method provides more detailed information, leveraging increasing computational resources. Moreover, the proposed method offers greater flexibility in setting thresholds and can be dynamically adjusted when the dataset changes, facilitating a more accurate determination of whether the detected lesion requires further biopsy treatment.
4.4 Generalizability Analysis on the ISIC Datasets
As illustrated in Table 4, we have predicted and compared the attributes present in the ISIC 2017 and 2018 datasets with the baseline utilized and the currently relevant SOTA, respectively. It is notable that our average AUC is the highest on both datasets, and specifically for each attribute, our proposed method attains the highest results in six out of nine attributes. This demonstrates that the proposed method is capable of effectively extracting the interactions between attributes, thereby enhancing the model’s predictive capability.
5 Discussion

### v3

@@ IV-C Evaluation Settings
AUC Inception-combined [6] 86.3 79.9 78.9 79.0 82.9 79.9 89.2 76.1 81.5
EmbeddingNet [44] 82.5 74.5 77.7 77.9 71.3 78.5 84.8 76.9 78.0
TripleNet [41] 81.2 73.8 76.3 77.6 76.8 76.0 85.1 79.9 78.3
Constrained-CC [25] 83.4 76.2 81.8 82.6 78.4 81.1 90.5 78.6 81.6
HcCNN [19] 85.6 78.3 77.6 81.3 81.9 82.6 89.8 82.7 82.5
GRM-DC [22] 87.6 87.5 81.2 83.6 79.0 83.1 90.8 75.4 83.5
CAFNet [30] 92.2 75.3 85.4 85.0 85.4 78.7 94.7 83.4 84.5
CCRD [42] 82.5 82.8 82.1 84.8 83.0 92.5 84.1 84.5
GELN [43] 90.2 84.0 82.3 83.8 82.9 81.9 91.8 84.5 85.2
RemixFormer++ [31] 91.3 87.7 81.6 82.5 86.1 84.2 91.1 86.2 86.3
Ours 92.0 86.7 90.1 88.3 89.3 86.2 93.4 83.0 88.6
Sens Inception-combined [6] 61.4 48.4 51.1 59.7 66.0 62.1 77.3 13.3 54.9
EmbeddingNet [44] 40.6 33.3 51.1 60.5 96.2 64.4 96.3 23.3 58.2
TripleNet [41] 46.5 33.3 39.4 61.3 97.9 67.2 90.0 30.0 58.2
Constrained-CC [25] 65.3 39.8 52.1 58.5 40.6 63.8 73.3 30.0 52.9
HcCNN [19] 58.4 40.9 35.1 55.7 95.2 80.2 92.2 20.0 59.7
GRM-DC [22] 59.0 77.5 67.0 39.2 21.9 70.1 69.9 3.6 51.0
CAFNet [30] 75.3 65.9 67.1 60.3 42.7 74.1 68.8 45.0 62.4
CCRD [42] 67.8 63.2 69.0 80.0 75.0 87.5 100.0 77.5
GELN [43] 68.1 57.4 57.0 68.9 74.6 70.0 76.7 50.0 65.3
RemixFormer++ [31] 72.3 61.3 67.0 63.7 47.2 76.8 72.0 33.3 61.7
Ours 74.3 63.4 76.6 56.5 79.2 85.9 82.7 40.0 69.8
Spec Inception-combined [6] 88.8 90.7 85.7 80.1 81.3 78.9 89.4 97.5 86.6
EmbeddingNet [44] 93.5 90.7 88.7 82.7 20.8 78.4 52.0 96.7 75.4
TripleNet [41] 90.1 93.0 92.4 76.8 23.6 71.6 60.0 98.6 75.8
Constrained-CC [25] 81.6 93.7 90.4 75.7 95.6 85.3 91.6 99.2 89.1
HcCNN [19] 88.1 92.4 90.0 86.3 41.5 71.6 65.3 98.4 79.2
GRM-DC [22] 89.5 79.0 80.3 95.8 96.8 78.8 91.0 100.0 88.9
CAFNet [30] 93.6 90.2 91.2 89.1 96.5 74.0 95.1 98.7 91.1
CCRD [42] 85.4 85.6 79.2 81.8 74.1 90.5 0.0 70.9
GELN [43] 87.7 83.5 84.5 81.9 81.6 76.8 91.3 92.6 85.0
RemixFormer++ [31] 93.9 89.7 79.7 84.5 94.8 75.2 89.7 97.3 88.1
Ours 90.8 89.1 85.4 87.5 81.7 65.6 89.1 95.8 85.6
V Results
V-A Comparison to State-of-the-Art Methods
Table I presents a detailed comparison of the classification performance between our proposed method and state-of-the-art deep learning-based techniques on the EDRA dataset. In this table, the top two AUC values for each label have been bolded to emphasize the highest-performing methods. Our approach achieved an average AUC of 88.6%, demonstrating strong generalization and robustness across multiple diagnostic labels. Specifically, it attained the highest AUC for four of the eight diagnostic labels: STR (90.1%), PIG (88.3%), RS (89.3%), and DaG (86.2%). For the remaining labels, it consistently ranked among the top two, underscoring its reliability across different classification tasks.
Building on this, when analyzing the comparative performance of existing models, all methods evaluated in Table I are deep learning-based. However, only GRM-DC and GELN incorporate GCNs, which explicitly model relationships between clinical attributes. Our approach further advances this direction by introducing the consideration of directed interactions between attributes and melanoma, leading to improved experimental results. These findings indicate that modeling directed relationships enhances diagnostic performance and contributes to a more interpretable model.
As the most recent state-of-the-art (SOTA) method aside from our proposed approach, RemixFormer++ integrates a transformer-based feature extractor with clinical diagnostic logic, achieving a strong performance. In contrast, our method further embeds clinical rules into the graph network while incorporating the self-attention mechanism used in transformers to enable multimodal task learning. This design reduces model size while enhancing performance. In the eight-label classification task, our approach achieved improvements in six categories, increasing the average AUC from 86.3% to 88.6%.
Although our proposed method achieves improvements in both AUC and sensitivity, its advantage over other methods in specificity (85.6%) is not as pronounced. This is primarily due to the detection mechanism of the 7PCL clinical algorithm, which is designed to maximize the identification of potential melanoma cases and ensure that positive patients are not overlooked. As a result, it inherently favors high sensitivity over specificity [45]. In our proposed method, we address this issue by optimizing the weight distribution and threshold selection, enabling the model to maintain a high sensitivity (69.8%) while achieving competitive specificity. This balanced approach enhances the reliability of the results and makes the model more clinically applicable.
V-B Ablation Studies
The proposed methodological framework consists of multiple modules, each serving a distinct role: CD-MFM facilitates comprehensive multimodal feature learning, 7PCL-DirGM enhances graph-based representation learning with directed interactions, and GD-DDW integrates clinical diagnostic knowledge to improve interpretability and decision-making. To systematically evaluate the contribution of each module, Table II presents a comparative analysis through an ablation study.
The first comparison assesses unimodal inputs, where dermoscopic and clinical images are processed separately, against multimodal fusion using dual-attention blocks. The results indicate that integrating dermoscopic and clinical imaging via dual-attention mechanisms improves the average AUC by 7% to 10%, highlighting the benefits of leveraging complementary information from different imaging modalities. Additionally, dermoscopic images alone outperform clinical images by an average of 3%, which aligns with expectations since dermoscopy provides enhanced feature visibility, particularly in the context of attribute-based analysis within the 7PCL framework.
Further analysis examines the individual impact of the 7PCL-DirGM and GD-DDW modules. When incorporated independently, both modules contribute to an average AUC improvement of 9% over the baseline dermoscopic model, with a particularly notable 13% increase in melanoma classification. These findings emphasize the critical role of incorporating structured graph-based relationships and gradient-based diagnostic decision weighting, both of which enhance the model’s ability to capture clinically relevant feature interactions.
The final evaluation considers the combined integration of all three modules, demonstrating a synergistic effect that further elevates model performance. The full framework achieves an average AUC of 87%, representing a significant enhancement over individual components. These results validate the effectiveness of combining multimodal feature learning, directed graph modeling, and diagnostic decision weighting, reinforcing the importance of clinically guided deep learning approaches for melanoma diagnosis.
@@ V-B Ablation Studies
TABLE II: Comparative Ablation Study of Key Modules in Our Proposed Method, Evaluated on AUC using the EDRA Dataset.
@@
Baseline (Clinical) Baseline (Derm) +Dual-Attention +7PCL-DirGM +GD-DDW Ours DIAG MEL 75.6 78.2 89.6 91.5 91.2 92.0 PN TYP 71.7 75.8 72.4 77.6 79.8 84.6 ATP 69.5 66.9 76.2 78.0 77.4 86.7 STR REG 79.2 84.2 79.1 81.2 82.7 88.1 IR 58.5 67.9 80.2 81.7 80.6 90.1 PIG REG 66.7 67.8 81.5 80.8 82.7 86.1 IR 68.1 72.0 81.2 81.0 81.1 88.3 RS PRS 58.4 72.9 76.2 78.6 78.9 89.3 DaG REG 69.6 67.7 80.0 81.8 82.2 84.5 IR 67.1 71.3 80.4 82.0 80.5 86.2 BWV PRS 79.6 83.6 92.3 93.3 92.3 93.4 VS REG 71.5 75.6 72.5 75.5 76.3 78.9 IR 66.4 64.3 73.6 76.6 75.8 83.0 Avg 69.4 72.9 79.6 81.5 81.7 87.0
@@
TABLE III: Comparison of Weights Value between Proposed Method and Traditional Algorithm under Different Tasks.
Task
@@
Mel vs Nevi Traditional 2 1 1 1 1 2 2
Proposed 1.81 1.36 1.20 0.87 0.93 2.35 1.66
Mel vs Others Traditional 2 1 1 1 1 2 2
Proposed 1.32 1.10 0.97 0.92 1.06 1.38 1.43
V-C Comparison to Traditional 7PCL Algorithm
The approach presented in this paper is primarily inspired by the 7PCL clinical scoring mechanism for melanoma diagnosis, with key advancements introduced through dynamic weight assignment and automated diagnosis enabled by GD-DDW. These improvements not only enhance diagnostic performance and expand the applicability of the model but also generate data-driven weighting information that can provide valuable feedback to frontline clinicians. By leveraging real-world data for adaptive weighting, the proposed method addresses the inherent subjectivity in predefined scoring systems and improves the model’s interpretability.
To systematically evaluate these improvements, we compare the performance of the proposed method with the traditional 7PCL approach in two distinct diagnostic scenarios. The first follows the conventional application of 7PCL, distinguishing between Mel and Nevi, where the original scoring system has been widely validated. The second considers a more complex diagnostic setting in which Mel must be differentiated from a broader range of skin diseases. This latter scenario reflects the challenges encountered in real-world clinical applications, where dermatological conditions exhibit significant variability. By assessing performance across these scenarios, we aim to highlight the benefits of adaptive learning and dynamic weight assignment in improving diagnostic performance beyond traditional rule-based approaches.
As shown in Table III, the traditional algorithm employs approximate categorization based on multivariate analysis [7], assigning weights of 2 or 1 to major and minor attributes, respectively. In contrast, the method proposed in this study automatically generates adaptive weights based on the prediction scores of each category under different tasks. Specifically, in the classification task of differentiating Mel from Nevi, the weights of the major attributes range from 1.7 to 2.4, while the weights of the minor attributes range from 0.8 to 1.4, closely resembling traditional algorithms. However, when the task shifts to distinguishing Mel from four other skin diseases, the weights dynamically adjust to range from 1.3 to 1.5 for major attributes and 0.9 to 1.1 for minor attributes.
This dynamic adjustment extends the applicability of the 7PCL algorithm, originally designed to distinguish Mel from Nevi, beyond its initial scope. As shown in Fig. 8b, the traditional algorithm experiences a significant drop in AUC, from 92% to 72%, when applied to more complex scenarios. In contrast, our proposed data-driven dynamic tuning approach maintains a high AUC value of 92% while using the same features in the classification tasks of Mel vs. Nevi and Mel vs. others. It should be noted that the data-driven weight parameters in Table III are derived from model training and can be dynamically optimized as more data becomes available. These results indicate that although these attributes were originally designed for distinguishing Mel from Nevi, their diagnostic relevance extends beyond this binary classification. By enabling automated dynamic weight adjustment, the proposed method effectively preserves the value of these attributes in differentiating Mel from a broader range of skin diseases. This adaptability highlights the potential of leveraging clinically established features while optimizing their contribution based on task-specific requirements, thereby improving the model’s generalization in more complex diagnostic scenarios.
V-D Generalizability Analysis on the ISIC Datasets
TABLE IV: Performance (AUC) of Proposed Method on ISIC Datasets.
@@ V-D Generalizability Analysis on the ISIC Datasets
ISIC2017 Milia-like cysts 63.2 68.3 74.7 75.4
Negative network 77.9 84.1 86.8 87.2
Pigment network 87.3 87.4 89.6 91.2
Streaks 96.2 93.9 95.6 95.2
Avg 81.2 83.4 86.6 87.3
ISIC2018 Globules 68.2 77.1 82.7 83.9
Milia-like cysts 60.1 68.8 74.3 76.4
Negative network 82.3 87.3 87.8 87.6
Pigment network 87.8 86.4 91.2 92.5
Streaks 84.3 88.7 91.8 91.6
Avg 76.5 81.7 85.6 86.4
Generalizability is a key factor in assessing a method’s clinical applicability. To evaluate this, we tested our approach on the ISIC 2017 and ISIC 2018 datasets and compared it with state-of-the-art methods. While these datasets contain only dermoscopic images, they encompass a diverse range of skin conditions, including melanoma, Nevi, and SK, and provide attribute annotations beyond the 7PCL framework, such as milia-like cysts. This evaluation assesses the model’s ability to generalize beyond 7PCL-specific attributes while accounting for dataset variations, demonstrating its robustness and adaptability to real-world dermatological diagnosis.
As shown in Table IV, we compared the predictive performance of our method against the baseline model (ResNet-50) and SOTA approaches (GRM-DC and RemixFormer++). Our method achieved the highest average AUC across both datasets, with 87.3% on ISIC 2017 and 86.4% on ISIC 2018. It also attained the highest classification performance in six out of nine attributes, further demonstrating its robustness. These results highlight the effectiveness of our method in extracting representative features and modeling attribute interactions, enhancing predictive performance for broader dermatological applications.
VI Discussion